Calculus I, Fall 2015 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 08

1. Section 4.2: problems 17, 18, 61, 65, 66.

2. Section 4.3: problems 7, 11, 74, 77, 80.

3. Section 4.4: problems 77, 85, 105, 111.